4 - (x) / (x-2) = (-2) / (x-2)
Show the steps that you would take to solve the following algebraically:

Answer by checkley77(12844)

4-(x)/(x-2)=(-2)/(x-2)
4-X/(X-2)+2/(X-2)=0
4-(X-2)/(X-2)=0
4-1=0
3=0